OpenStudy (anonymous):

Are these ratios equivalent? 3/5 and 9/15

OpenStudy (fibonaccichick666):

can you simplify 9/15 at all?

sammixboo (sammixboo):

Take a look at this website! It explains lots about Equivalent Ratios! http://www.virtualnerd.com/tutorials/?id=PreAlg_06_02_0001

sammixboo (sammixboo):

It is saying equivalent ratios are just like equivalent fractions, so if two ratios have the same value, then they are equivalent. All you need to do to find out if ratios are equivalent is to simplify all the ratios, and if the ratios you simplified are the same, then they are equivalent!

OpenStudy (anonymous):

find a common denominator and compare

sammixboo (sammixboo):

So can \(\rm \large \color{red}{\frac{3}{5}}\) can be simplified any further? Can \(\rm \large \color{red}{\frac{9}{15}}\) be simplified any further?

OpenStudy (anonymous):

Yes, because 9/3=3 and 15/3=5